### v3

# Sim-to-real supervised domain adaptation for radioisotope identification

Abstract
Machine learning has the potential to improve the speed and reliability of radioisotope identification using gamma spectroscopy. However, meticulously labeling an experimental dataset for training is often prohibitively expensive, while training models purely on synthetic data is risky due to the domain gap between simulated and experimental measurements. In this research, we demonstrate that supervised domain adaptation can substantially improve the performance of radioisotope identification models by transferring knowledge between synthetic and experimental data domains. We consider two domain adaptation scenarios: (1) a simulation-to-simulation adaptation, where we perform multi-label proportion estimation using simulated high-purity germanium detectors, and (2) a simulation-to-experimental adaptation, where we perform multi-class, single-label classification using measured spectra from handheld lanthanum bromide (LaBr3) and sodium iodide (NaI(Tl)) detectors. We begin by pretraining a spectral classifier on synthetic data using a custom transformer-based neural network. After subsequent fine-tuning on just 64 labeled experimental spectra, we achieve a test accuracy of 96% in the sim-to-real scenario with a LaBr3 detector, far surpassing a synthetic-only baseline model (75%) and a model trained from scratch (80%) on the same 64 spectra. Furthermore, we demonstrate that domain-adapted models learn more human-interpretable features than experiment-only baseline models. Overall, our results highlight the potential for supervised domain adaptation techniques to bridge the sim-to-real gap in radioisotope identification, enabling the development of accurate and explainable classifiers even in real-world scenarios where access to experimental data is limited.
keywords:
Domain Adaptation, Transformer, Neural Network, Gamma Spectroscopy1 Introduction
Spectroscopy and spectrometry are widely used characterization methods, and depending on the spectrum type, can illuminate details about a material’s crystallographic, molecular, elemental, or isotopic composition. The applications of such spectral characterization and analysis span fields including high-energy physics [1], remote sensing [2], biology [3], and nuclear material analysis [4]. While the motivations, physical mechanisms, and equipment used vary widely across spectral analysis, they are all unified by the processing of a “spectrum”–the histogram of discrete counts of a quantum arriving at a detector versus a physical quantity (usually energy or wavelength) indicative of each quantum over time. In the present work, we focus on gamma spectroscopy, which measures a spectrum of the number of gamma photons depositing energy in a detector across discrete ranges of deposited energy [5]. While we focus on specifics of gamma spectroscopy, the analysis of how to best enable spectral simulation-based inference is more general.
Many detector types have been developed for detecting gammas and their energy, each with unique resolution (a measurement of the ability to discriminate between small differences in gamma energy), efficiency (a measurement of the proportion of gammas incident on the detector that are detected), and calibration (a description of how the deposition energy is converted to an electrical signal and eventually measured). Traditional techniques for gamma spectral analysis search for peaks in the measured gamma energy spectrum that can then be mapped to specific decay energies of a radioisotope [6, 7, 8]. These methods typically include numerous time-consuming preprocessing steps and require close monitoring by an expert spectroscopist [9, 10, 11]. Furthermore, classical approaches struggle in the presence of thick shielding or when the spectrum contains many overlapping material signatures. Supervised machine learning is an excellent alternative candidate for spectral analysis due to its ability to learn near-optimal representations of data for a given task [12]. Gamma spectroscopy, however, introduces some challenges to the adoption of supervised machine learning. First and foremost, supervised machine learning requires large labeled datasets for high performance, and so-called “neural scaling laws” have emerged showing enormous scale is required for difficult tasks (e.g., 38 million images per class for 99% accurate image classification on natural images [13]). The collection of such large datasets of gamma spectroscopy data would be extremely costly and time-consuming. Further, the labeling of a gamma spectrum is ambiguous. At the spectrum level, if a complete spectrum is labeled according to a recorded set of experimenter-placed radioactive sources, these labels will be incomplete due to naturally occurring radioactive material (NORM) in the background and cosmic radiation sources; at the quanta level, it is presently impossible to label each detection event according to its origin source from among the nearby radioactive sources, NORM, and cosmic sources [14].
This has led to a trend in the field of supervised learning for spectroscopic analysis where simulation-based inference is leveraged [14, 15, 16, 17]. Amongst simulation-based inference techniques, naive application to real data of a classifier trained on simulated data incurs a huge misclassification risk, since the ability of such models to generalize is inherently limited by the difference in the data distributions between those generated by simulation and those generated by real processes, hereafter the “sim-to-real gap”. In this analysis, we consider the scenario where we have access to a data-rich source domain (simulations) and wish to transfer knowledge to perform classification in a data-scarce target domain (experimental measurements). In the present work, we characterize the effectiveness of this approach, leaving the investigation of other simulation-based inference approaches, such as the iterative approaches from Ref. [15] or unsupervised and semi-supervised approaches [18] for future studies. As such, we consider two primary research questions:
1.
Does pretraining followed by fine-tuning yield a statistically significant improvement in model performance over models trained (a) only on source-domain data and (b) only on target-domain data?
2.
Do different architectures produce statistical differences in the chosen evaluation metrics?
Furthermore, we are interested in how the answer to these questions depends on (i) target domain dataset size, (ii) simulated versus experimental target domain, and (iii) detector type (HPGe, LaBr3, NaI(Tl)).
2 Related Work
Machine learning methods were first used for radioisotope identification by Olmos et al. [19, 20], where the authors demonstrated that artificial neural networks (ANNs) can achieve high classification accuracy on experimental gamma spectra. More recently, Liang et al. trained a convolutional neural network (CNN) using simulated mixed sources, achieving high accuracy on low-resolution spectra [21]. Kamuda developed an open-source Python package annsa, capable of creating simulated gamma-ray training datasets and applying machine learning models to solve spectroscopic tasks [22]. The author built upon this research in Ref. [23], finding that CNNs had smaller variance in their output than MLPs.
Moore et al. applied the CNN architecture to time-sequenced gamma spectra by analyzing a two-dimensional waterfall plot [24]. The authors continued this research in Ref. [25], where they achieved mixed results leveraging transfer learning techniques to combine simulated data with experimental measurements. Daniel et al. also considered the task of out-of-distribution (OOD) testing, achieving accuracy on real data after training a CNN on a fully synthetic database [26].
Bilton et al. examined the use of ANNs for mobile detection using simulated NaI(Tl)detectors, finding that autoencoder-based spectral anomaly detection outperforms PCA and NMF by up to and recurrent neural network enhanced source identification boosts accuracy by over static ANNs [27]. Ghawaly et al. similarly trained a deep convolutional autoencoder, outperforming principal component analysis (PCA) based methods at flagging anomalous gamma spectra in both real-world and synthetic urban environments [28].
Khatiwada et al. surveyed different machine learning algorithms for emergency response applications. The authors found that models trained on synthetic datasets offered excellent performance on simulated test sets, and performance was comparable to conventional methods when applied to an experimental dataset [29]. Li et al. introduced a more advanced transformer-based neural network for nuclide identification, achieving an improved recognition rate compared to classical architectures [30]. The authors reshaped 1024-channel gamma spectra into patches, and the z-score-normalized counts of each patch were used as the embedding vector for the subsequent attention layers.
Morrow et al. developed an open-source Python package PyRIID for generating synthetic gamma spectra for training machine learning models [31]. Van Omen et al. used PyRIID to synthesize short-lived fission products to train a model using a semi-supervised loss function, aiming to perform OOD multi-label proportion estimation [32]. Barradas et al. similarly used high-quality synthetic spectra to train a complex ANN, outperforming conventional peak-recognition methods when identifying up to ten radionuclides in multi-source LaBr3 gamma spectra [33].
3 Dataset Curation
We consider six datasets, divided into three domain adaptation scenarios, summarized in Table 1. In the first scenario (sim-to-sim adaptation), the task is multi-label proportion estimation on a simulated dataset of mixed geant spectra using a HPGe detector. In the second and third scenarios (sim-to-real adaptation), the task is multi-class, single-label classification on a dataset of experimental handheld spectra taken using LaBr3 and NaI(Tl) detectors.
| Scenario | Domain |
Size | ||||||
| sim-to-sim (HPGe) | Source | gadras | Mixed | 55 | |||||||
| Target | geant | Mixed | 55 | ||||||||
| sim-to-real (LaBr3) | Source | gadras | Single | 32 | |||||||
| Target | Experiment | Single | 32 | 15,091 | |||||||
| sim-to-real (NaI(Tl)) | Source | gadras | Single | 32 | |||||||
| Target | Experiment | Single | 32 | 10,440 |
3.1 Sim-to-sim adaptation
In the first scenario, we simulated a dataset of gamma spectra using gadras (source domain) and aim to perform isotopic identification on a spectral dataset simulated in geant (target domain). In geant, we simulated the gamma emission spectra of 55 individual radioisotopes by uniformly dissolving each isotope in a scintillation cocktail. A high-purity germanium (HPGe) detector was placed one meter from the source, and the energy of all incoming gamma rays was recorded and binned to create a gamma spectrum. The simulations were repeated using a mix of different scintillation cocktails and HPGe detector models in order to capture additional variance which may be present in a realistic environment. We leveraged the G4ARES simulation framework to improve estimations of gamma cascade summing corrections [16]. In post-processing, we applied a low-energy cutoff and performed Gaussian energy broadening over each spectrum [5]. Furthermore, we introduced slight variations to energy calibration and detector resolution between different detector setups. Once these template spectra were simulated, we created mixed spectra by arbitrarily summing up to 14 individual spectra with random proportions. Our choice to include multiple isotopes per spectrum captures the common scenario in which several radioisotopes are simultaneously present in a material sample. We subsequently added background (cosmic, 40K, 226Ra, and 232Th) and Poisson noise. We synthesized a total of training spectra, along with separate holdout sets for validation and testing. To clarify, we did not use the entire dataset for model training; instead, we trained on data subsets of varying sizes to compare the performance of different models (with and without pretraining) as a function of dataset size.
We followed a similar sequence of steps for generating gadras data, first simulating 55 template spectra using a 95% efficiency HPGe detector. The spectra were then mixed, background-added, and Poisson resampled to produce a total of training spectra, performed using the PyRIID SeedMixer and StaticSynthesizer classes [31]. In both gadras and geant, all spectra consisted of 1024 uniform energy bins spanning from 0 to 3000 keV, with slight endpoint fluctuations due to variations in energy calibration.
3.2 Sim-to-real adaptation
In the experimental domain adaptation scenarios, we again use gadras as our source domain, but seek to perform isotopic classification using experimental spectra. The experimental dataset was collected in 2018 at Pacific Northwest National Laboratory to characterize the performance of handheld LaBr3 and NaI(Tl) detectors. 32 different source isotopes were used, with either steel, lead, polyethylene, or no shielding. The data were preprocessed to remove spectra that were corrupted, empty, or missing metadata, as well as to interpolate all spectra onto a uniform 1024-bin energy grid from 0 to 3000 keV. This resulted in a labeled dataset of 15,091 LaBr3 spectra and 10,440 NaI(Tl) spectra, partitioned into training, validation, and test sets with a 70/15/15 split. As before, we only considered subsets of the training data at a time to benchmark performance as a function of training size. It is worth mentioning that the experimental datasets available for this study are relatively simple, and 100% testing accuracy can be achieved while only training on training examples due to the nature of the data collection process (single-label, high signal-to-noise ratio). While a more complicated experimental dataset is desirable, we still expect the results of this research to apply broadly and refer to the sim-to-sim adaptation as a case study using complex, mixed sources.
The gadras dataset was synthesized in an attempt to best replicate the experimental configuration. Spectra were simulated for each of the 32 radioisotopes over a range of source-to-detector distances, detector heights, energy calibrations, and shielding configurations. These spectra were then mixed with gadras-simulated background (cosmic, 40K, 226Ra, and 232Th) to create seed templates using the PyRIID SeedMixer class. The mixing fractions were drawn from a Dirichlet distribution whose parameters were estimated by fitting measured background spectra as a non-negative combination of the four simulated component spectra. These mixed spectrum seed templates were then used to create a synthetic dataset using the PyRIID StaticSynthesizer class, varying over parameters such as background count rate, live time, and signal-to-noise ratio, with the values of each parameter approximated empirically. In total, we synthesized training spectra, validation spectra, and testing spectra in less than 15 minutes on a personal laptop. This process was repeated for each detector type (LaBr3, NaI(Tl)).
4 Method
4.1 Model Architectures
We implemented four machine learning architectures for this analysis. Multilayer perceptrons (MLPs) and convolutional neural networks (CNNs) are commonly used in the gamma-spectroscopy literature and serve as simple, effective benchmarks. We also consider attention-based models (first introduced for neural machine translation [34]) and include a prior transformer-based neural network for nuclide identification, TBNN (Li) [30]. Finally, we introduce TBNN (ours), a more general transformer that modifies the Vision Transformer (ViT) [35] for source identification in gamma spectroscopy (Fig. 1). Below, we motivate each architecture class by the assumptions it encodes about spectral data and the implications for generalization:
- Multilayer Perceptron
MLPs use one or more fully connected layers that project the input via learned weight matrices, apply a nonlinear activation function, and pass the result to subsequent layers. MLPs can be universal function approximators, and therefore have extremely high capacity for learning classification rules from data. Unfortunately, there are no inductive biases within MLPs, and they must encounter data examples of all realistic features. In the gamma spectroscopy domain, this manifests itself as MLPs struggling to generalize across unseen calibrations and resolutions.
- Convolutional Neural Networks
CNNs incorporate one-dimensional convolutional layer(s) that slide learnable filters across local segments of the spectrum. To overcome issues with MLPs in a finite lattice domain (i.e., in spectra, where channels are equally spaced along one dimension, or images where pixels are equally spaced along two dimensions), CNNs assume that information from channels close to each other is correlated, and that the same feature will appear the same way in different spatial areas of the lattice. In gamma spectroscopy, this should enable generalization to shifts in calibration, although the a priori choice of convolution size may be difficult to tune for detectors of different resolutions.
- Transformers
For data representable as sequences, many different sequence models have been proposed and tested, progressing from recurrent neural networks (RNNs) [12], Transformers [34], and recent selective state space models (SSMs) [36]. Transformers are the most common choice due to their ability to train in a fully parallel manner, and the amount of literature available characterizing them. In gamma spectroscopy, their attention mechanism should be able to determine higher order interactions between the number of counts in different channels, which is analogous to what gamma spectroscopists perform implicitly. Limitations in interaction range are often cited for Transformers, so widely separated peaks from a specific isotope may be difficult to capture.
In TBNN (Li) [30], the authors first reshape a 1024-dimensional spectral input into a sequence and then add fixed sinusoidal positional encodings. The resulting embeddings are then fed into a stack of multi-head self-attention blocks, with a final dense layer mapping to the predicted isotopes. We introduce a more general transformer-based neural network, TBNN (ours), featuring several modifications over Li et al. [30]:
1.
Learnable patch embeddings via a small CNN to map counts in each patch into embedding vectors, rather than using the counts directly. A simple linear layer or MLP could also be used.
2.
Learnable [CLS] token to aggregate information from all patches to produce the final prediction rather than simple flattening.
3.
Tunable patch size and embedding dimension to optimize the trade-off between local detail preservation and computational efficiency.
4.
Pre-norm residual ordering to improve training stability and convergence.
5.
Learnable positional encoding to allow the model to adapt the positional information based on the specific characteristics of the spectral data.
1.
4.2 Training Protocols
For each domain adaptation scenario, we trained three classes of models using the datasets generated in Section 3. The first class, source-only, was a single model simply trained on the entire source domain (gadras) dataset. The second class of models, target-only, were trained solely on the target-domain dataset (geant in the sim-to-sim scenario; experiment in the sim-to-real scenario). Within this class, we trained multiple different models, each using a random subset of the target domain data of size for , where is the total number of training examples. The third class of models, domain-adapted, were pretrained using the source-domain dataset, and subsequently fine-tuned using the target-domain data. As before, multiple domain-adapted models were trained with data subsets of varying sizes to evaluate the effect of dataset size on model performance. For all models, input spectra were first z-score normalized, which is a common data preprocessing technique in machine learning to stabilize training [37].
4.3 Hyperparameter Search
Prior to training, we performed an in-depth Bayesian hyperparameter search using the Optuna software package [38]. We determined architecture hyperparameters using the source domain dataset in the sim-to-sim scenario with validation loss as our search criterion. To ensure a direct comparison, these architecture hyperparameters were used for all models in this study, while training hyperparameters (learning rate, batch size, dropout, weight decay) were calculated separately for each domain adaptation scenario. Since we might expect the optimal training hyperparameters to vary with dataset size, we performed a separate hyperparameter search for every data subset size. The fine-tuning hyperparameters were determined separately from the target-only hyperparameters, and the fine-tuning runs considered which (if any) layers to freeze as part of the hyperparameter search. Similar to the results of Ref. [39], we sometimes found that freezing the output layer(s) while fine-tuning only the earlier layers yielded better performance. We provide a further discussion of the hyperparameter tuning process in Section A.2.
4.4 Uncertainty Estimation
We employed an ensemble method to estimate the uncertainty in our model performance. Each model was trained with 10 different random initial configurations, and the variations in performance were used to quantify model uncertainty. For the source-only model class, this simply meant training 10 models for each architecture using different random weight initializations and seed configurations. For the target-only and domain-adapted model classes, we additionally sampled a random subset of target-domain data during each random trial. These trials were paired by ensuring that the random data subset was identical between target-only and domain-adapted models with the same seed. Similarly, each fine-tuning trial was initialized from the source-only pretraining checkpoint corresponding to its seed, so the measured variance reflects both pretraining and fine-tuning.
5 Results
5.1 Predictive Performance (Classification Metrics)
After training three classes of models (source-only, target-only, domain-adapted) for each of the three domain adaptation scenarios (sim-to-sim HPGe, sim-to-real LaBr3, sim-to-real NaI(Tl)) with different architectural backbones (MLP, CNN, TBNN (Li), TBNN (ours)) for a range of different dataset sizes, we recorded the performance of each model on a held-out target-domain testing dataset. The choice of a best metric to be used for spectroscopic analysis has no consensus in the literature, although the detection limit and quantification uncertainty are preferred by spectroscopists for the detection and quantification of individual isotopes [4, 14, 40]. Without selecting details about the amount of background present and the isotopes of interest, these metrics are not computable, and traditional machine learning metrics characterizing the classifiers themselves must be used111Unfortunately, some evidence in [14] shows that accuracy may not be monotonic with detection limit or quantification accuracy, so caution should be used when extending classifier metrics to real-world performance.. For the sim-to-real scenarios, we used testing accuracy as our evaluation metric. For the sim-to-sim scenario, accuracy is no longer a well-defined metric since we are performing multi-label proportion estimation; instead, we computed the testing absolute proportion error (APE) score as our evaluation metric. The APE score (Eq. 1) provides a convenient 0-to-1 scoring metric to compare how precisely different models predict the fractional composition of each mixed source. We provide a more detailed discussion of the APE score as an evaluation metric in Section A.1.
5.1.1 Sim-to-sim comparison
We summarize the results of the sim-to-sim scenario in Fig. 3, where we compare the geant APE score across different model classes for each architecture. Overall, we observe that the APE scores of the domain-adapted models very often match or exceed those of the source-only and target-only models. In particular, using a fine-tuning dataset of 1024 spectra, the TBNN (ours) model achieves a testing APE score of , outperforming both the source-only model (APE score = ) and the target-only model (APE score = ) on the same dataset size. We quantify these results in Table 2(a) where we explicitly record the testing APE scores for models trained with 0 versus 64 versus 1024 target domain training spectra.
To test the statistical significance of these conclusions, we performed a series of one-sided Wilcoxon signed-rank tests using the output of 10 paired trials with random initial configurations, as described in Section 4.4. In these tests, the null hypothesis states that the difference in APE scores between the domain-adapted models and the baseline models (source-only or target-only) is nonpositive, and we use a significance threshold of . These results are indicated in the bottom panels of Fig. 3, where we can reject the null hypothesis for nearly all cases, except domain-adapted versus source-only at small fine-tuning sizes.
5.1.2 Sim-to-real comparison
We perform the same analysis for the sim-to-real comparisons and summarize the results in Figs 4 and 5, comparing the testing accuracy across different model classes and architectures. These results are quantified in Tables 2(b) and 2(c). Similar to the sim-to-sim scenario, we observe a statistically significant improvement in testing accuracy for most training sizes. In particular, using 64 fine-tuning spectra, the TBNN (ours) model achieves an accuracy of on the LaBr3 testing dataset, easily surpassing the source-only model (accuracy = ) and the target-only model (accuracy = ) trained on the same experimental subset. We see a similar result on the NaI(Tl) dataset, achieving an accuracy of using a TBNN (ours) architecture and 64 fine-tuning spectra, versus of the source-only model and of the target-only model.
| Target-only | Source-only | Domain-adapted | |||
| N=64 | N=1024 | N=0 | N=64 | N=1024 | |
| MLP | |||||
| CNN | |||||
| TBNN (Li) | |||||
| TBNN (ours) | |||||
| Target-only | Source-only | Domain-adapted | |||
| N=64 | N=1024 | N=0 | N=64 | N=1024 | |
| MLP | |||||
| CNN | |||||
| TBNN (Li) | |||||
| TBNN (ours) | |||||
| Target-only | Source-only | Domain-adapted | |||
| N=64 | N=1024 | N=0 | N=64 | N=1024 | |
| MLP | |||||
| CNN | |||||
| TBNN (Li) | |||||
| TBNN (ours) | |||||
5.1.3 Transformer vs. Classical Architectures
To quantify architectural differences in model performance, we computed each model’s testing evaluation metric for all fine-tuning dataset sizes. For each dataset size, we performed two-sided Wilcoxon signed-rank tests to test the null hypothesis that there is no difference in testing score between each model architecture. In the sim-to-sim scenario (Fig. 6(a)), the transformer architectures (TBNN (Li) and TBNN (ours)) perform strongly across most sizes: TBNN (ours) is statistically superior at small sizes (-), while TBNN (Li) is best at intermediate sizes (-). The MLP only leads at large sizes (), and the CNN does not lead at any size. We quantify these results in Table 7(a), where overall, domain-adapted transformer architectures demonstrate a clear statistical advantage over classical architectures across a wide range of fine-tuning sizes. Notably, these trends do not appear in the target-only model class, suggesting that transformers benefit more from pretraining than classical architectures. We hypothesize that this result reflects the higher model complexity and greater data requirements of transformers, and that rigorous synthetic pretraining can partially mitigate this large-data hurdle.
In the sim-to-real scenarios (Figs. 6(c) and 6(b)), we do not observe a statistically significant difference between model architectures for most dataset sizes. As indicated in Tables 7(b) and 7(c), there is not a strong, consistent trend supporting one model architecture over the others. We hypothesize that the relative simplicity of the sim-to-real classification problem offers minimal advantages to using more complex architectures, and thus recommend that future work perform similar architectural comparisons on more challenging experimental classification scenarios.
5.2 Model Analysis (Diagnostics & Explainability)
In Section 5.1, we present statistical evidence that fine-tuning yields an improvement to model performance as measured by classification metrics (accuracy, APE score). However, a comparison between the accuracy of two models offers little insight as to why one model outperforms the other, or whether such improvement would necessarily transfer to a new research scenario. To enrich this analysis, we include comparisons across several diagnostic metrics (calibration, separation, and smoothness), as well as qualitative understanding using SHapley Additive exPlanations (SHAP).
5.2.1 Model Diagnostics
In addition to classification accuracy, we consider a range of well-established metrics from literature that quantify calibration, separation, uncertainty, and input-space smoothness. Calibration quantifies a model’s confidence in its predictions, penalizing the gap between a model’s predicted confidence and its actual accuracy. Separation evaluates how decisively a model makes its predictions by comparing the confidence in the chosen class to the next most likely class. Uncertainty measures a model’s confusion between different classes, penalizing scenarios where a model’s prediction is split among several classes. Smoothness assesses the model’s decision boundary, rewarding models that make similar predictions for similar inputs. Our goal with these comparisons is to include grounded, complementary techniques to offer insights into model performance beyond pure classification accuracy.
We present our results in Table 3 using a TBNN (ours) architecture for the sim-to-real LaBr3 adaptation scenario. In addition to standard classification accuracy (acc), we include negative log-likelihood (nll), which more heavily penalizes overconfidence [41]. The Brier score (brier) measures the mean-squared error between the true one-hot encoded label and predicted probability vectors [42], while the expected calibration error (ece) bins predictions by confidence to directly measure the confidence-accuracy gap [43]. Prediction margin evaluates the mean (margin_mean) and 10th percentile (margin_p10) of the margin for each spectrum, defined as the difference in predicted probabilities between the top two classes [44]. Mean predictive entropy (entropy_mean) measures how confident or uncertain the model is in its predictions by calculating the average Shannon entropy of the predicted probability vectors [45]. Mean Jacobian norm (jacobian_norm_mean) quantifies a model’s sensitivity to its input by averaging the squared -norm of the gradient of the loss with respect to the input across all samples [46]. Lastly, we considered four boundary complexity and smoothness metrics by first building a -NN graph using a Euclidean distance metric and on the input -normalized spectra of the testing dataset. We considered a range of distance metrics (Euclidean, cosine, correlation) and different values for (5-30), finding minimal changes in results. We then considered the post-softmax probability vectors for each edge and computed the following metrics [47, 48]:
1.
knn_tv_hard: fraction of pairs with different predicted hard labels (i.e., how often ).
2.
knn_prob_l2: mean squared distance between probability vectors (i.e., ).
3.
knn_conf_absdiff: average absolute change in confidence between probability vectors (i.e., ).
4.
knn_margin_absdiff: average absolute change in true-class margin (i.e., define , where is the ground-truth class for sample . Then, take ).
To test the statistical significance of each diagnostic metric, we ran a one-sided Wilcoxon signed-rank test comparing the empirically best model to the second best model across 10 paired random trials. The results of this section are summarized in Table 3, where we see that the source-only model wins in 1 out of 12 metrics, target-only wins in 2 out of 12 metrics, and domain-adapted in 7 out of 12 metrics, with 2 out of 12 metrics considered statistically insignificant. In particular, the domain-adapted model performs well in the calibration, separation, and sensitivity metrics, suggesting more trustworthy predictions. The input-space smoothness metrics yield mixed results, suggesting that fine-tuning primarily improves reliability as opposed to dramatically altering the input-space geometry of the decision boundary.
Metric Source-only Target-only 64 Domain-adapted 64 -value (best vs. 2nd) Dir. Type acc Accuracy nll [41] Likelihood / Loss brier [42] Calibration ece [43] Calibration margin_mean [44] Separation / Margin margin_p10 [44] Separation / Margin entropy_mean [45] Uncertainty jacobian_norm_mean [46] Sensitivity knn_tv_hard [47, 48] Boundary Complexity knn_prob_l2 [47, 48] Smoothness knn_conf_absdiff [47, 48] Smoothness knn_margin_absdiff [47, 48] Smoothness
5.2.2 Model Interpretability
We also present evidence that beyond metric scores, domain-adapted models learn more human-interpretable features compared to target-only models. In Explainable AI literature, various techniques have been proposed for quantifying input feature relevance, such as gradient-based [49, 50] and Attention Rollout [51] methods. In Ref. [52], the authors considered various explanation methods in the context of gamma spectroscopy, and concluded that SHAP offered especially accurate explanations of radioisotope predictions compared to other methods. In short, SHAP evaluates an input feature’s importance by removing the feature and assessing the resulting effect on the model’s output. A more mathematically thorough description of the KernelSHAP algorithm can be found in Refs. [52, 53].
For this analysis, we chose a target-only and domain-adapted TBNN (ours) model from the sim-to-real LaBr3 adaptation scenario, each trained with the same 64 experimental spectra. Fig. 2 shows a 99Mo spectrum overlaid with a color plot of calculated SHAP explanation values for each region. Large, positive SHAP values (highlighted in red) indicate that a feature positively correlates with the model’s class prediction. The SHAP explanation reveals that the target-only model relies exclusively on the 140.5 keV peak from 99mTc when computing its classification, ignoring the 739.5 keV peak characteristic of 99Mo. As a result, the target-only model frequently confuses 99mTc and 99Mo due to reliance on the shared daughter line, resulting in an overall false alarm rate of 2.0% for 99Mo. Conversely, the domain-adapted model correctly identifies both the 140.5 keV and 739.5 keV peaks, enabling unambiguous discrimination between both isotopes.
Our analysis further reveals that target-only models will often overfit to the low energy region of the spectrum. As seen in Fig. 7, when classifying 239Pu, the target-only model identifies the - keV band as the most salient component of the spectrum, indicative of overfitting. Conversely, the domain-adapted model’s SHAP explanation most strongly highlights a region containing the 345, 375, and 413 keV peaks, which are commonly used by spectroscopists to identify 239Pu, especially in the presence of shielding [54]. Overall, we observe that domain-adapted models more consistently produce explanations that align with human-interpretable photopeaks, while target-only models are more likely to focus on low-energy, high-count features. As such, we hypothesize that gadras pretraining confers an interpretability boost, even when the target-only and domain-adapted models have similar per-class accuracy.
6 Conclusion
In this study, we present a supervised domain adaptation framework for fine-tuning models pretrained on synthetic data to perform classification in an experimental downstream setting. We consider both sim-to-sim and sim-to-real adaptation scenarios, and find that domain-adapted models provide a statistically significant improvement in classification metrics compared to synthetic-only models or experiment-only models with no pretraining. For example, using a custom transformer-based neural network implementation, we achieve a testing accuracy of 96% after fine-tuning on 64 LaBr3 spectra, improved from 75% of a source-only model and 80% using a target-only model. Furthermore, we find that domain-adapted models identify more human-interpretable features compared to target-only models. These results highlight a path forward to apply machine learning techniques to experimental settings where only a limited amount of experimental training data is available.
7 Acknowledgements
This research was supported by the Laboratory Directed Research and Development Program at Pacific Northwest National Laboratory, a multiprogram national laboratory operated by Battelle for the U.S. Department of Energy under contract DE-AC05-76RLO1830. Peter Lalor is grateful for the support of the Linus Pauling Distinguished Postdoctoral Fellowship. The authors would like to acknowledge Tyler Morrow and Brian Archambault for their useful suggestions and feedback. The authors declare no conflict of interest.
Appendix A Supplementary Materials
A.1 Defining an evaluation metric for mixed proportion estimation
Choosing an evaluation metric for multi-label proportion estimation is less intuitive than for traditional classification problems. In multi-label proportion estimation scenarios, several isotopes may be present simultaneously, with the fractions of all isotopes summing to unity, rendering a simple metric such as accuracy elusive. Instead, we define the absolute proportion error (APE) score, computed by rescaling the mean absolute error to yield a value between 0 and 1. Explicitly,
In Eq. 1, is the number of spectra in the testing dataset, is the number of radioisotope classes, and and are matrices of the predicted and true proportion labels, respectively. An APE score of 1 indicates perfect agreement between the predicted and true isotope proportions, while a score of 0 reflects maximal deviation. This evaluation metric provides a convenient way to compare how precisely different models predict the fractional composition of each mixed source.
A.2 Hyperparameter search
For each architecture (MLP, CNN, TBNN (Li), TBNN (ours)), we conducted a Bayesian hyperparameter search with 150 trials on the sim-to-sim source domain dataset. Each model was trained for 20 minutes on an NVIDIA Tesla V100 GPU with 32 GB of memory, with the number of epochs adjusted dynamically to maintain a consistent overall training time. Underperforming configurations were pruned early, and an early stopping criterion was applied to mitigate overfitting. Hyperparameters were ranked by their validation cross-entropy loss, and the best-performing parameters are shown in Table 4. Source-only baseline models for all scenarios and all architectures were trained for 80 minutes on the same hardware, and the training was repeated 10 times using different random seeds.
When training the target-only and domain-adapted models, the same architecture hyperparameters were used as before, but training hyperparameters (learning rate, batch size, dropout, weight decay) were recalculated using a subsequent Bayesian hyperparameter search with 50 trials and 5 minutes per trial. The fine-tuning runs additionally included which layers to freeze (if any) as part of the hyperparameter search. We conducted a separate hyperparameter search for each scenario as well as for every data subset size. Table 5 lists the best hyperparameters for data subsets of size 64 and 1024 for the target-only and domain-adapted runs for each of the three domain adaptation scenarios.
| Architecture | Parameter | Search Space | Best run |
| MLP | Learning Rate | {1e-5, 2e-3} | 5.12e-5 |
| Batch Size | {32, 512} | 512 | |
| Weight Decay | {1e-7, 1e-1} | 2.67e-5 | |
| Dropout | {0.0, 0.4} | 0.346 | |
| Num Dense Layers | {1, 4} | 2 | |
| Dense1 Hidden Units | {512, 8192} | 4096 | |
| Dense2 Hidden Units | {256, 4096} | 2048 | |
| CNN | Learning Rate | {1e-5, 2e-3} | 1.17e-4 |
| Batch Size | {32, 512} | 512 | |
| Weight Decay | {1e-7, 1e-1} | 3.42e-5 | |
| Dropout | {0.0, 0.4} | 0.101 | |
| Num Convolutional Layers | {1, 3} | 1 | |
| Conv Filters | {16, 256} | 32 | |
| Conv Kernel Size | {3, 9} | 7 | |
| Num Dense Layers | {1, 2} | 2 | |
| Dense1 Hidden Units | {512, 8192} | 2048 | |
| Dense2 Hidden Units | {256, 4096} | 1024 | |
| TBNN (Li) | Learning Rate | {1e-5, 2e-3} | 1.02e-3 |
| Batch Size | {32, 512} | 256 | |
| Weight Decay | {1e-7, 1e-1} | 3.46e-3 | |
| Dropout | {0.0, 0.4} | 4.53e-4 | |
| Num Attention Blocks | {1, 8} | 5 | |
| Num Heads | {1, 8} | 4 | |
| FF Dimension | {64, 8192} | 1024 | |
| TBNN (ours) | Learning Rate | {1e-5, 2e-3} | 1.49e-4 |
| Batch Size | {32, 512} | 64 | |
| Weight Decay | {1e-7, 1e-1} | 5.02e-3 | |
| Dropout | {0.0, 0.4} | 0.0198 | |
| Embedding Method | {Linear, MLP, CNN} | CNN | |
| CNN Embed Filters | {8, 256} | 8 | |
| Embedding Dimension | {8, 1024} | 256 | |
| Num Attention Blocks | {1, 5} | 4 | |
| Num Heads | {1, 8} | 8 | |
| FF Dimension | {2, 4096} | 512 | |
| Patch Size | {16, 64} | 64 | |
| Positional Encoding | {Sinusoidal, Learnable} | Learnable |
(a) Training hyperparameters for the sim-to-sim gadrasgeant adaptation using a HPGe detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 9.58e-5 9.01e-4 2.00e-4 3.82e-4 Batch Size {32, 512} 32 512 64 512 Weight Decay {1e-7, 1e-1} 4.08e-4 2.92e-3 3.17e-2 3.69e-5 Dropout {0.0, 0.4} 0.373 0.337 0.397 0.395 Frozen Layers {none, all} N/A N/A None None CNN Learning Rate {1e-6, 1e-3} 1.92e-5 7.16e-4 9.68e-4 8.71e-5 Batch Size {32, 512} 32 512 64 64 Weight Decay {1e-7, 1e-1} 7.76e-2 2.84e-4 9.84e-7 5.36e-7 Dropout {0.0, 0.4} 0.381 0.275 0.303 0.322 Frozen Layers {none, all} N/A N/A Last 3 layers First 2 layers TBNN (Li) Learning Rate {1e-6, 1e-3} 6.73e-4 9.75e-4 6.44e-4 6.53e-4 Batch Size {32, 512} 64 128 32 512 Weight Decay {1e-7, 1e-1} 2.73e-7 2.14e-2 1.71e-5 6.86e-5 Dropout {0.0, 0.4} 0.149 0.067 0.027 0.021 Frozen Layers {none, all} N/A N/A None None TBNN (ours) Learning Rate {1e-6, 1e-3} 3.37e-4 7.67e-4 1.82e-4 5.77e-5 Batch Size {32, 512} 32 32 64 512 Weight Decay {1e-7, 1e-1} 1.31e-2 4.43e-5 5.48e-4 1.36e-6 Dropout {0.0, 0.4} 0.051 0.159 0.175 0.186 Frozen Layers {none, all} N/A N/A None Last 5 layers (b) Training hyperparameters for the sim-to-real gadrasexperiment adaptation using a LaBr3 detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 1.19e-6 1.44e-6 8.10e-6 2.78e-4 Batch Size {32, 512} 64 32 64 256 Weight Decay {1e-7, 1e-1} 1.45e-6 8.26e-2 4.24e-7 3.80e-5 Dropout {0.0, 0.4} 0.021 0.098 0.000312 0.344 Frozen Layers {none, all} N/A N/A Last 2 layers Last layer CNN Learning Rate {1e-6, 1e-3} 2.69e-6 5.06e-5 1.13e-4 4.49e-4 Batch Size {32, 512} 64 32 32 256 Weight Decay {1e-7, 1e-1} 9.62e-5 3.58e-2 1.06e-5 2.33e-5 Dropout {0.0, 0.4} 0.000239 0.0349 0.00476 0.389 Frozen Layers {none, all} N/A N/A First 3 layers None TBNN (Li) Learning Rate {1e-6, 1e-3} 2.65e-5 8.79e-4 5.07e-4 8.85e-4 Batch Size {32, 512} 32 128 32 256 Weight Decay {1e-7, 1e-1} 2.62e-4 6.98e-5 1.10e-2 1.30e-4 Dropout {0.0, 0.4} 0.0117 0.0125 0.156 0.0125 Frozen Layers {none, all} N/A N/A Last layer Last 8 layers TBNN (ours) Learning Rate {1e-6, 1e-3} 6.95e-6 2.25e-4 7.02e-4 2.37e-4 Batch Size {32, 512} 32 32 32 128 Weight Decay {1e-7, 1e-1} 1.63e-4 2.76e-5 4.54e-7 1.49e-4 Dropout {0.0, 0.4} 0.000792 0.0347 0.0476 0.370 Frozen Layers {none, all} N/A N/A Last 19 layers Last layer (c) Training hyperparameters for the sim-to-real gadrasexperiment adaptation using a NaI(Tl) detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 1.01e-6 7.79e-6 2.37e-4 4.84e-4 Batch Size {32, 512} 32 32 64 32 Weight Decay {1e-7, 1e-1} 4.05e-7 1.78e-2 1.49e-4 4.03e-4 Dropout {0.0, 0.4} 0.012 0.019 0.370 0.141 Frozen Layers {none, all} N/A N/A None Last layer CNN Learning Rate {1e-6, 1e-3} 2.14e-6 1.66e-4 7.43e-6 2.27e-4 Batch Size {32, 512} 64 256 64 32 Weight Decay {1e-7, 1e-1} 2.32e-3 4.70e-3 4.61e-5 4.51e-5 Dropout {0.0, 0.4} 0.00335 0.00206 0.0120 0.00117 Frozen Layers {none, all} N/A N/A Last layer None TBNN (Li) Learning Rate {1e-6, 1e-3} 6.39e-4 3.75e-4 1.84e-4 7.74e-4 Batch Size {32, 512} 64 64 64 32 Weight Decay {1e-7, 1e-1} 1.30e-4 4.65e-4 2.33e-4 7.06e-4 Dropout {0.0, 0.4} 0.00170 0.0298 0.00506 0.0383 Frozen Layers {none, all} N/A N/A First 5 layers First 13 layers TBNN (ours) Learning Rate {1e-6, 1e-3} 9.29e-5 2.35e-4 7.31e-4 2.53e-4 Batch Size {32, 512} 64 32 32 64 Weight Decay {1e-7, 1e-1} 2.13e-2 6.41e-7 4.98e-2 6.62e-5 Dropout {0.0, 0.4} 0.00192 0.0164 0.184 0.350 Frozen Layers {none, all} N/A N/A First 3 layers None
MLP CNN TBNN (Li) TBNN (ours) Target Dataset Size Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only to
MLP CNN TBNN (Li) TBNN (ours) Target Dataset Size Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only
MLP CNN TBNN (Li) TBNN (ours) Target Dataset Size Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only Domain-adapted vs. source-only Domain-adapted vs. target-only
| Fine-tuning Dataset Size | MLP vs. Best | CNN vs. Best | TBNN (Li) vs. Best | TBNN (ours) vs. Best |
| 0 | 0.002 | 0.002 | 0.002 | BEST |
| 0.002 | 0.002 | 0.002 | BEST | |
| 0.002 | 0.002 | 0.002 | BEST | |
| 0.002 | 0.002 | 0.002 | BEST | |
| 0.002 | 0.002 | 0.002 | BEST | |
| 0.002 | 0.002 | 0.004 | BEST | |
| 0.002 | 0.002 | 0.002 | BEST | |
| 0.002 | 0.002 | BEST | 0.492 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.002 | 0.002 | BEST | 0.010 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.002 | 0.002 | BEST | 0.002 | |
| 0.232 | 0.002 | BEST | 0.002 | |
| BEST | 0.002 | BEST | 0.037 | |
| BEST | 0.002 | 0.002 | 0.002 | |
| BEST | 0.002 | 0.002 | 0.002 |
Fine-tuning Dataset Size MLP vs. Best CNN vs. Best TBNN (Li) vs. Best TBNN (ours) vs. Best 0 0.002 0.002 BEST 0.002 BEST 0.027 0.084 BEST 0.232 0.020 BEST 0.084 0.002 BEST 0.027 0.625 0.002 0.004 BEST 0.415 0.432 0.049 BEST 0.010 BEST 0.027 0.221 0.006 BEST 0.041 BEST 0.838 0.011 0.014 0.441 BEST 0.076 0.011 0.153 BEST 0.036 0.074 0.305 BEST 0.579 0.579 BEST BEST BEST BEST BEST BEST BEST BEST BEST BEST
Fine-tuning Dataset Size MLP vs. Best CNN vs. Best TBNN (Li) vs. Best TBNN (ours) vs. Best 0 0.002 BEST 0.492 0.010 0.002 0.002 BEST 0.006 0.027 0.002 BEST 0.064 0.275 0.002 BEST 0.160 0.105 0.002 BEST 0.027 0.193 0.432 BEST 0.432 0.010 BEST 0.695 0.160 0.002 0.153 0.281 BEST 0.006 0.021 0.560 BEST 0.305 BEST BEST 0.579 BEST BEST BEST BEST BEST BEST BEST BEST BEST BEST BEST BEST